C-504 Adam and Forrest have wrapped up their season and are back in McMurdo.
C-506 The stream team completed their stream monitoring this week. They closed out the gage boxes for the winter and did end-of-season lake levels. Aneliya is working on Chlorophyll analysis for Algae Ops. The team is unpacking and sorting all of the retro. They have 2 more day trips planned for this week.
C-507 finished their work and went north earlier this week.
C-509 spent a day sampling the Kaki Ponds, aka Marr Ponds, see photo attached. The analytical team has been making good progress on lab work, but then had a series of issues with all of the analytical instruments. As of 1/30/16 none of the instruments were working properly. As of now, anion analyses are working, cations are still unknown and both Shimadzu instruments are having significant problems and we continue to work on troubleshooting. We are working to resolve all of these issues and making contingency plans as we wrap up the season.